Hello,

I need help using a system of equations to solve this problem:

Trinidad has $2.35 in nickels and dimes. If she has 33 coins in all, find the number of nickels and dimes.

Thank you

Answer by scott8148(5895)

let x=nickels y=dimes ... x+y=33 ... x=33-y

5x+10y=235 ... 5(33-y)+10y=235 ... 165-5y+10y=235 ... 5y=70 ... y=14

x=33-14 ... x=19